OpenStudy (julyahx1):

Are the following figures similar? Parallelograms ABCD and EFGH are shown. AB equals 2. AD equals 10. EF equals 5. EH equals 25. Yes; the corresponding angles are congruent Yes; the corresponding sides are proportional No; the corresponding angles are not congruent No; the corresponding sides are not proportional I'm thinking it's C, but i'm not so sure. Please someone help me. Thank you!

OpenStudy (julyahx1):

@563blackghost please help me

OpenStudy (julyahx1):

@3mar

OpenStudy (3mar):

With my pleasure!

OpenStudy (julyahx1):

@TTP

OpenStudy (julyahx1):

Thank you!

OpenStudy (3mar):

Can you free-hand sketch them here?

OpenStudy (julyahx1):

thats a pic of it

OpenStudy (3mar):

Nice presentation! If you notice: \[\frac{ EH }{ AD }=\frac{ FE }{ BA }=\frac{ 25 }{ 10 }=\frac{ 5 }{ 2 }=2.5\] The sides are proportional!

OpenStudy (julyahx1):

wait

OpenStudy (julyahx1):

it is?

OpenStudy (julyahx1):

B?

OpenStudy (julyahx1):

@3mar so it is B not C?

OpenStudy (3mar):

B I think as we have been given about sides not angles!

OpenStudy (julyahx1):

but are you 100% sure